Assume that the following events occurred at a division of Generic Electric for March of the current year:

  

1. Purchased $100 million in direct materials.

2. Incurred direct labor costs of $57 million.

3. Determined that manufacturing overhead was $83 million.

4. Transferred 80 percent of the materials purchased to work-in-process.

5. Completed work on 75 percent of the work-in-process. Costs are assigned equally across all work-in-process.

6. The inventory accounts have no beginning balances. All costs incurred were debited to the appropriate account and credited to Accounts Payable.

Required:

Give the amounts for the following items in the Work-in-process account: (Do not round your intermediate calculations. Enter your final answers in millions rounded to 2 decimal places.)
Answer #1

Answer:

(a) Transfers-in (TI) = 57+83+(100*80%)= $220 million
(b) Transfers-out (TO) = 220*75%= $165 million
(c) Ending balance (EB) = 220-165= $55 million